### v2

# Cycle Matroids of Graphings: From Convergence to Duality

Abstract
A recent line of research has concentrated on exploring the links between analytic and combinatorial theories of submodularity, uncovering several key connections between them. In this context, Lovász initiated the study of matroids from an analytic point of view and introduced the cycle matroid of a graphing [18]. Motivated by the limit theory of graphs, the authors introduced a form of right-convergence, called quotient-convergence, for a sequence of submodular setfunctions, leading to a notion of convergence for matroids through their rank functions [3].
In this paper, we study the connection between local-global convergence of graphs and quotient-convergence of their cycle matroids. We characterize the exposed points of the associated convex set, forming an analytic counterpart of matroid base-polytopes. Finally, we consider dual planar graphings and show that the cycle matroid of one is the cocycle matroid of its dual if and only if the underlying graphings are hyperfinite.
Keywords: Cycle matroid, Duality, Exposed points, Graphings, Matroid limits, Cost
1 Introduction
Cycle matroids bridge graph theory and matroid theory by providing matroid theoretic counterparts of graph concepts such as cycles, cuts, and spanning trees, thus leading to a powerful framework for understanding and solving various combinatorial and optimization problems. The matroidal perspective allows for a deeper analysis of graph properties and leads to efficient algorithms in network design, electrical engineering, and operations research.
Graphings are bounded degree graphs on standard probability measure spaces, serving as limit objects in sparse graph limit theory, see Subsection 2.2 for the definition. Recently, Lovász [18] initiated the study of cycle matroids from an analytic point of view and introduced the cycle matroid of a graphing $\mathcal{G}=(X,E,\mu)$ by defining its rank function as
$\rho_{\mathcal{G}}(F)=1-\mathbb{E}_{x}\left[\frac{1}{|V(\mathcal{G}^{F}_{x})|}\right],$ (1)
where $F\subseteq E$ is Borel, $x$ is a $\mu$-random vertex of $\mathcal{G}$ and $\mathcal{G}^{F}_{x}$ denotes its connected component in the graphing $\mathcal{G}^{F}=(X,F,\mu)$. It was shown that $\rho_{\mathcal{G}}$ is an increasing submodular setfunction on the Borel subsets of $E$. It is not difficult to verify that this definition generalizes the normalized rank of the cycle matroid of a finite graph.
This raises several questions about cycle matroids of graphings. How to define and describe their bases, cycles, and other basic properties known from the finite case? Does the convergence of a sequence of finite graphs to a graphing imply convergence of the corresponding matroids (in the sense introduced in [3])? Does duality of matroids play any role? In this paper, we provide some answers to these questions. We develop the limit theory of these matroids, study their bases, and investigate how they behave under duality of planar graphings. We give some background on earlier work on cycle matroids of infinite graphs in Subsection 1.5.
1.1 Limit theory
Our first result connects local-global convergence of graphs with quotient-convergence of the rank function of cycle matroids, defined in [3]. We give a short definition, but refer to Subsection 2.4 for the setup and a precise formulation. A $k$-quotient of a setfunction $\varphi$ on a set-algebra $(J,\mathcal{B})$ is $\varphi\circ F^{-1}$, where $F:J\to[k]$ is a measurable map. We denote by $Q_{k}(\varphi)$ the set of $k$-quotients of $\varphi$. A sequence of setfunctions $(\varphi_{1},\varphi_{2},\dots)$ quotient converges to $\varphi$, denoted $\varphi_{n}\rightarrowtail\varphi$, if $Q_{k}(\varphi)\to Q_{k}(\varphi)$ in the Hausdorff distance for all $k\in\mathbb{N}$.
Theorem 1.1.
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to the graphing $\mathcal{G}$ in the local-global sense. Let $r_{G_{n}}$ denote the rank function of the cycle matroid of $G_{n}$. Then
$\frac{r_{G_{n}}}{|V(G_{n})|}\rightarrowtail\rho_{\mathcal{G}}$
in the sense of quotient-convergence of submodular setfunctions.
We also exhibit an example showing that assuming local convergence of $G_{n}$ is not enough.
Theorem 1.2.
There exists a sequence $(H_{1},H_{2},\dots)$ of finite graphs with all degrees bounded by $D$ that converges to a graphing $\mathcal{G}$ in the local sense, however
[…]
1.2 Fractional bases
For a finite matroid $M=(E,r)$ with ground set $E$ and rank function $r$, the set of fractional bases
$C_{M}=\left\{x\in\mathbb{R}^{E}~{}\Big{|}~{}x\geq 0,\ \sum_{e\in A}x(a)\leq r(A)\text{ for }A\subseteq E,\ \sum_{e\in E}x(e)=r(E)\right\}$
forms an integral polytope whose vertices are exactly the bases of $M$, called the base-polytope of $M$. For the cycle matroid of a connected graph, this means that among fractional spanning trees, the extreme ones are genuine spanning trees.
In the analytic setting developed by Lovász [19], the analogue of a base-polytope is the set of basic minorizing finitely additive measures. In particular, for a graphing $\mathcal{G}=(X,E,\mu)$, this gives
$C_{\mathcal{G}}=\text{bmm}^{+}(\rho_{\mathcal{G}})=\big{\{}\alpha\textrm{ finitely additive measure on $E$}~{}\big{|}~{}\alpha\leq\rho_{\mathcal{G}},\ \alpha(E)=\rho_{\mathcal{G}}(E)\big{\}}.$
A vertex of a convex polytope generalizes to the infinite setting in several ways. A point $v$ of the convex set $C$ is extreme if it cannot be written as $v=\frac{1}{2}(u+w)$, where $u,w\in C\backslash\{v\}$. The point $v$ is exposed if there is a supporting hyperplane $H$ through $v$ such that $H\cap C=\{v\}$. This property depends on the topology of the ambient vector space; in this paper, we mean that $H$ is defined by the inner product with a function in $L^{1}(E)$. An exposed point is necessarily extreme, but not the other way around; see Section 4 for further comments.
An essential spanning forest of a graphing $\mathcal{G}$ is an acyclic subset $T$ of the edges that forms a spanning tree in finite components of $\mathcal{G}$, and contains only infinite $T$-components in the infinite $\mathcal{G}$-components. Note that $T$ might contain several components in an infinite component of $\mathcal{G}$, but all of those are required to be infinite; this is why $T$ is called a spanning forest instead of a spanning tree. As a generalization of the finite case, we prove the following.
Theorem 1.3.
Let $\mathcal{G}$ be a graphing. Then, the exposed points of $C_{\mathcal{G}}$ are exactly the hyperfinite essential spanning forests of $\mathcal{G}$.
The following question remains open.
Question 1.4.
Are all extreme points of $C_{\mathcal{G}}$ exposed, and therefore hyperfinite spanning forests?
1.3 Matroids beyond the hyperfinite
While the definition of the cycle matroid via (1) is a consistent generalization of the finite notion, it does have a drawback: the rank of the full edge set is at most 1. This is natural for finite graphs, as spanning trees have average degree slightly below 2. For graphings, however, spanning trees can have higher average degree. The notion of cost is designed to measure the minimal amount of edges needed to generate the connected components of a graphing; see the precise definition in Subsection 2.5. It seems natural to ask if one can come up with a notion of cycle matroid whose total rank is equal to the cost.
Question 1.5.
Is there a notion of the cycle matroid of a graphing $\mathcal{G}$ with submodular rank function $\varphi_{\mathcal{G}}$ such that $\varphi_{\mathcal{G}}(E)=\textrm{cost}(\mathcal{G})$?
A positive answer to this question would not invalidate $\rho_{\mathcal{G}}$, rather complement it. See e.g. Theorems 1.9 and 1.10 below, where we use ideas from [7] to give a positive answer for a certain class of planar graphings.
To further justify the study of $\rho_{\mathcal{G}}$, we highlight that even though $\rho_{\mathcal{G}}$ only sees spanning forests that are hyperfinite, it still carries a lot of information about the graphing. In the finite case, Whitney proved that every 3-connected graph is uniquely determined by its cycle matroid. In an upcoming work [2], we study analogous problems for graphings.
Question 1.6.
Under what assumptions does $\rho_{\mathcal{G}}$ determine the graphing $\mathcal{G}$ (up to measure zero)?
As a first step in this direction, we could show that $\rho_{\mathcal{G}}$ already determines acyclic graphings with minimum degree 3 up to isomorphism in the following sense: For a graphing $\mathcal{G}$ and an acyclic graphing $\mathcal{T}$ of minimum degree 3, if $\rho_{\mathcal{G}}\cong\rho_{\mathcal{T}}$ then $\mathcal{G}\cong\mathcal{T}$. In both cases, $\cong$ means that there exists a measure preserving bijection between the ground spaces, up to measure zero, that maps one structure to the other.
1.4 Duality
Duality is an important concept of matroid theory, and it lines up nicely with duality of finite planar graphs: the dual of a cycle matroid of an (embedded) planar graph is the cycle matroid of its planar dual. The question naturally arises: Does an analogous statement hold for planar graphings?
The notion of a planar graphing takes some work to set up. We postpone this to Section 5, where we give an informal introduction and refer to [1] and [7] for a rigorous treatment. For now, we only say that in a planar graphing, the connected components are planar maps, i.e. they are endowed with a specified planar embedding. We assume that these planar maps have no accumulation points in the plane. We also want to point out that the notion of a “dual graph” is not completely evident when infinite faces are present [24, 25].
When we talk about a pair of dual graphings $\mathcal{G}$ and $\mathcal{G}^{*}$, it is assumed, informally speaking, that the components of $\mathcal{G}^{*}$ are, as planar maps, the duals of the components of $\mathcal{G}$. For simplicity, we exclude infinite faces, we assume that components are $2$-vertex-connected to avoid edges that are only adjacent to one face, and that the maximum size of the faces is bounded to make sure that both $\mathcal{G}$ and $\mathcal{G}^{*}$ have bounded degree. For a dual pair, there is a Borel bijection $\sigma:~{}E(\mathcal{G})\to E(\mathcal{G}^{*})$ such that the image of the star of a vertex of $\mathcal{G}$ is a facial cycle in $\mathcal{G}^{*}$ and vice versa. We also note that the hyperfiniteness of $\mathcal{G}$ and $\mathcal{G}^{*}$ are equivalent [1, Example 3.11, Proposition 3.12].
Remark 1.7.
The above assumptions are made solely to avoid technical difficulties, not because they are necessary for the results presented here. Our aim in this paper is to showcase how duality works, and these assumptions allow us to do so with minimal distraction. The aforementioned works [7] and [1] demonstrate that, with dedicated treatment, such technical questions can be addressed to prove theorems for planar graphings without these restrictions.
The dual of the cycle matroid, called the cocyclic matroid of $\mathcal{G}$, can be defined for any (not necessarily planar) graphing by its rank function
$\rho^{D}_{\mathcal{G}}(A)=\rho_{\mathcal{G}}(E\setminus A)+\tilde{\mu}(A)-\rho_{\mathcal{G}}(E),$
where $\tilde{\mu}$ denotes the edge measure of $\mathcal{G}$. Our first result here shows that in the hyperfinite case duality works as in the finite case.
Theorem 1.8.
Let $\mathcal{G}$ and $\mathcal{G}^{*}$ be a hyperfinite pair of dual planar graphings. Then, $\rho^{D}_{\mathcal{G}}=\rho_{\mathcal{G}^{*}}$ and $\rho^{D}_{\mathcal{G}^{*}}=\rho_{\mathcal{G}}$.
However, when $\mathcal{G}$ and $\mathcal{G}^{*}$ are not hyperfinite, the two matroids do not coincide. This is the case for example if the components of $\mathcal{G}$ and $\mathcal{G}^{*}$ are dual hyperbolic tilings. We can actually exploit this phenomenon to give a partial positive answer to Question 1.5.
Theorem 1.9.
Let $\mathcal{G}$ and $\mathcal{G}^{*}$ be a non-hyperfinite pair of dual planar graphings. Then, $\rho^{D}_{\mathcal{G}^{*}}\neq\rho_{\mathcal{G}}$. Moreover, $\rho^{D}_{\mathcal{G}^{*}}(E)$ is a submodular function on $E(\mathcal{G})$ that attains the cost, i.e.
$\rho^{D}_{\mathcal{G}^{*}}(E)=\textrm{cost}(\mathcal{G}).$
Note that planar graphings are treeable [7], so if they are not hyperfinite, their cost is strictly more than $1$. Both submodular functions $\rho_{\mathcal{G}}$ and $\rho^{D}_{\mathcal{G}^{*}}$ are natural to consider, as supported by the following result.
Theorem 1.10.
Let $\mathcal{G}$ and $\mathcal{G}^{*}$ be a pair of dual planar graphings. Let $<$ be an arbitrary Borel total order on the edges. The Wired Maximal Spanning Forest $\mathcal{T}^{<}_{w}$ is an exposed point of $\text{bmm}^{+}(\rho_{\mathcal{G}})$, while the Free Maximal Spanning Forest $\mathcal{T}^{<}_{f}$ is an exposed point of $\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$.
In [1, Theorem 1 (15)], it was shown that a unimodular random map is hyperfinite if and only if the wired and free maximal spanning forests coincide. In our setup, this can be interpreted as $\mathcal{T}^{<}_{f}=\mathcal{T}^{<}_{w}$ in the special case when $\mathcal{G}$ is an (edge) Bernoulli graphing over the unimodular random planar map, and $<$ is the natural ordering on the edges given by the edge-labeling.
1.5 Other notions of cycle matroids on infinite graphs
In this paper we study the cycle matroids of graphings, as graphings are natural limit objects in the context of sparse graph limits. It is equally natural to extend the notion of cycle matroid to infinite graphs without a measurable structure, for example connected locally finite graphs. There has been an extensive line of research on this topic. Although it is not directly related to our results, we mention some of the highlights.
Probably the simplest notion for infinite graphs is the finitary cycle matroid whose circuits are exactly the finite cycles. However, an unfortunate feature of finitary matroids is that excluding infinite circuits spoils duality. As a workaround, Bruhn, Diestel, Kriesell, Pendavingh and Wollan [6] proposed five equivalent sets of axioms leading to a non-finitary theory of infinite matroids. These axioms were in fact shown to be equivalent to the notion of B-matroids introduced by Higgs [11]. Using these axioms, Bruhn and Diestel [5] provided further examples by defining the circuits of the matroid to be the edge sets of topological circuits of compactifications of infinite graphs obtained by adding the space of ends to the graph. Later, Bowler, Carmesin and Christian [4] introduced a class of infinite cycle matroids that unifies previous constructions.
2 Preliminaries
2.1 Submodular setfunctions and matroids
A setfunction $\varphi\colon\mathcal{B}\to\mathbb{R}$ on a set-algebra $(J,\mathcal{B})$ is submodular if $\varphi(X)+\varphi(Y)\geq\varphi(X\cap Y)+\varphi(X\cup Y)$ holds for $X,Y\in\mathcal{B}$. One of the most basic examples of a submodular function is the rank function of a matroid. A matroid $M=(E,r)$ is defined by its finite ground set $E$ and its rank function $r\colon 2^{E}\to\mathbb{Z}_{+}$ that satisfies the so-called rank axioms: (R1) $r(\emptyset)=0$, (R2) $X\subseteq Y\Rightarrow r(X)\leq r(Y)$, (R3) $r(X)\leq|X|$, and (R4) $r(X)+r(Y)\geq r(X\cap Y)+r(X\cup Y)$. Note that axiom (R4) requires submodularity of the rank function. A set $X\subseteq E$ is called independent if $r(X)=|X|$ and dependent otherwise. The circuits of the matroid are formed by its inclusionwise dependent sets, while maximal independent sets are called bases. The rank axioms imply that any two bases have the same cardinality. For further details on matroids, the reader is referred to [22].
2.2 Graphings and hyperfiniteness
We give the basic definitions for completeness, but point to [17, Chapters 18 and 21] for a thorough introduction. Let $(X,\mu)$ be a Borel probability space. A graphing is a graph $\mathcal{G}$ with $V(\mathcal{G})=X$ and Borel edge set $E(\mathcal{G})\subseteq X\times X$, in which all degrees are finite, and
[…]
for all measurable sets $A,B\subseteq X$, where $\deg_{S}(x)$ is the number of edges from $x\in X$ to $S\subseteq X$. The assumption (2) allows one to meaningfully define the edge measure $\tilde{\mu}$ on $X\times X$ by setting
[…]
on product sets, and extending by Caratheodory’s theorem. The measure $\tilde{\mu}$ is concentrated on $E(\mathcal{G})$. A graphing $\mathcal{G}$ is hyperfinite if, for every $\varepsilon>0$, there is a Borel set $S\subseteq E(\mathcal{G})$ with $\tilde{\mu}(S)\leq\varepsilon$ such that every connected component of $E\setminus S$ is finite. An acyclic graphing is called a treeing. A treeing is hyperfinite exactly when its components have 00, $1$, or $2$ ends.
2.3 Local and local-global convergence
Intuitively, a sequence $(G_{1},G_{2},\dots)$ of finite graphs is said to be locally (or Benjamini-Schramm) convergent if the local behaviour around a random vertex of $G_{n}$ converges as $n\to\infty$. Formally, let $r\in\mathbb{N}$ denote our radius of sight, and $\alpha$ denote a rooted, connected graph of radius $r$. The local statistics $P_{r}(G,\alpha)$ of a finite graph $G$ are the probabilities $\mathbb{P}_{v}\big{[}B_{G}(v,r)\cong\alpha\big{]}$, where the vertex $v\in V(G)$ is chosen uniformly at random, $B_{G}(v,r)$ denotes the $r$-ball around $v$ in $G$, and the isomorphism has to respect the root. One can define $P_{r}(G,\alpha)$ analogously for a graphing $\mathcal{G}$, with the random vertex being chosen according to $\mu$. A sequence $(G_{1},G_{2},\dots)$ of finite graphs converges locally to a graphing $\mathcal{G}$ if the local statistics converge, i.e. $P_{r}(G_{n},\alpha)\to P_{r}(\mathcal{G},\alpha)$ for all $r$ and $\alpha$.
Local-global convergence is a more refined notion of convergence, see [10] for an introduction. We use a colored version of the neighborhood statistics. For a fixed vertex-coloring $c\colon V(G)\to[k]$, the colored neighborhood statistics $P_{r}(G,c,\alpha)$ is obtained as before: we choose a vertex $v\in V(G)$ uniformly at random, and then consider its colored $r$-neighborhood $\big{(}B_{G}(v,r),c|_{B_{G}(r,v)}\big{)}$. Then the probabilities $P_{r}[G,c]=\big{(}P_{r}(G,c,\alpha)\big{)}_{\alpha}$ form a distribution on the set of rooted, colored, connected graphs $\alpha$ of radius at most $r$. Let $Q_{k,r}(G)$ denote the (finite) set of possible colored neighborhood statistics we can obtain by different colorings:
$Q_{k,r}(G)=\big{\{}P_{r}[G,c]~{}\big{|}~{}c\colon V(G)\to[k]\big{\}}.$
The set $Q_{k,r}(\mathcal{G})$ is defined analogously for a graphing, with the assumption that the coloring $c\colon V(\mathcal{G})\to[k]$ is required to be Borel. The sequence $G_{n}$ locally-globally converges to $\mathcal{G}$, if $Q_{k,r}(G_{n})\to Q_{k,r}(\mathcal{G})$ in the Hausdorff distance. The local-global convergence of $G_{n}$ means that for any $r,k\in\mathbb{N}$ and $\varepsilon>0$, there is a threshold such that, for $i$ large enough, the colored neighborhood distribution $P_{r}[G_{i},c]$ for any $k$-coloring $c$ can be approximately modeled on $\mathcal{G}$. That is, we can find some Borel coloring $c^{\prime}$ of $\mathcal{G}$ such that $P_{r}[G_{i},c]$ and $P_{r}[\mathcal{G},c^{\prime}]$ are $\varepsilon$-close. Vice versa, Borel colorings of $\mathcal{G}$ can be approximately modeled on $G_{i}$.
2.4 Quotient-convergence of setfunctions
The quotient-convergence of setfunctions was introduced in [3]. Let $\varphi\colon\mathcal{B}\to\mathbb{R}$ be a setfunction on a set-algebra $(J,\mathcal{B})$ with $\varphi(\emptyset)=0$. The $k$-quotient set $Q_{k}(\varphi)$ is then defined as the set of all quotients of $\varphi$ on $[k]$, that is, the set of all setfunctions $\varphi\circ F^{-1}$ where $F\colon J\to[k]$ is a measurable map. Note that $Q_{k}(\varphi)$ is a subset of $\mathbb{R}^{2^{k}}$, with the two coordinates $\varphi\circ F^{-1}(\emptyset)=0$ and $\varphi\circ F^{-1}(J)=\varphi(J)$ being fixed. The quotient profile of $\varphi$ is defined as the sequence $(Q_{1}(\varphi),Q_{2}(\varphi),\dots)$, while the sequence of closures $(\overline{Q}_{1}(\varphi),\overline{Q}_{2}(\varphi),\dots)$ is the closed quotient profile of $\varphi$. A sequence $(\varphi_{1},\varphi_{2},\dots)$ of setfunctions on set-algebras $(J_{1},\mathcal{B}_{1}),(J_{2},\mathcal{B}_{2}),\dots$ is quotient-convergent if for every positive integer $k$, the quotient sets $(Q_{k}(\varphi_{1}),Q_{k}(\varphi_{2}),\dots)$ form a Cauchy sequence in the Hausdorff distance $d_{{\text{\rm Haus}}_{2^{k}}}$ on the $2^{k}$-dimensional Euclidean space. The sequence $(\varphi_{1},\varphi_{2},\dots)$ quotient-converges to $\varphi$ for some setfunction $\varphi$ if the sequence $(Q_{k}(\varphi_{1}),Q_{k}(\varphi_{2}),\dots)$ converges to $Q_{k}(\varphi)$ in the Hausdorff distance, denoted by $\varphi_{n}\rightarrowtail\varphi$. For a finite measurable partition $\mathcal{P}=\{P_{1},\dots,P_{q}\}$ of $J$, that is, each $P_{i}$ is measurable, then $\varphi/\mathcal{P}\colon 2^{[k]}\to\mathbb{R}$ denotes the map defined by $(\varphi/\mathcal{P})(X)=\varphi(\cup_{i\in X}P_{i})$ for $X\subseteq[k]$.
2.5 Cost
We give a very brief introduction to the theory of cost, and refer the reader to [8] and [12] for more details. For two graphings $\mathcal{G}$ and $\mathcal{H}$ on the same vertex set $(X,\mu)$ we write $\mathcal{G}\sim\mathcal{H}$ if they have the same connected components. The cost of $\mathcal{G}$ is defined to be
$\textrm{cost}(\mathcal{G})=\frac{1}{2}\inf\big{\{}\tilde{\mu}_{\mathcal{H}}\big{(}E(\mathcal{H})\big{)}\ \big{|}\ \mathcal{H}\sim\mathcal{G}\big{\}}.$
Gaboriau proved that one cannot “cut costs” on treeings, that is, their edge measure is the same as their cost.
Theorem 2.1 (Gaboriau [9]).
For a treeing $\mathcal{T}=(X,E,\mu)$, we have $\textrm{cost}(\mathcal{T})=\mu_{\mathcal{T}}(E)$.
For a graphing with all components of size $n$, its cost is $1-\frac{1}{n}$, and for hyperfinite graphings with infinite components the cost is 1. These facts were observed by Levitt, and we combine them to fit our setup in the following theorem.
Proposition 2.2 (Levitt  [15]).
Let $\mathcal{T}=(X,E,\mu)$ be a treeing. Then, the following are equivalent:
  1. (i)
    $\mathcal{T}$ is hyperfinite,
  2. (ii)
    $\textrm{cost}(\mathcal{T})=\rho_{\mathcal{T}}(E)$.
3 Connection to sparse graph limits
In this section we prove Theorems 1.1 and 1.2.
3.1 Continuity for local-global convergence
It was proved already in [18, Theorem 4.7] that the rank of the whole edge set of a graph is locally estimable, i.e. continuous with respect to local convergence [17, Theorem 22.3].
Theorem 3.1.
Let $(G_{1},G_{2},\dots)$ be a sequence of finite graphs with all degrees bounded by $D$ that converges to a graphing $\mathcal{G}$ in the local sense. Then, $r_{n}(E(G_{n}))\to\rho_{\mathcal{G}}(E(\mathcal{G}))$.
We can strengthen this result to imply Theorem 1.1 by the following observation. Given any graphing parameter $f$, we can turn it into a setfunction $\varphi_{f}$ on subsets of edges of $\mathcal{G}=(X,E,\mu)$ by
$\varphi_{\mathcal{G},f}(F)=f\big{(}(X,F,\mu)\big{)}.$
Theorem 3.2.
If $f$ is continuous with respect to local convergence and $\mathcal{G}_{n}\to\mathcal{G}$ in the local-global sense, then $\varphi_{\mathcal{G}_{n},f}\rightarrowtail\varphi_{\mathcal{G},f}$.
Proof.
For sake of brevity, let $\varphi_{n}=\varphi_{\mathcal{G}_{n},f}$, $\varphi=\varphi_{\mathcal{G},f}$, $Q^{n}_{k}=Q_{k}(\varphi_{n})$ and $Q_{k}=Q_{k}(\varphi)$. We need to prove that for all $k$, $Q^{n}_{k}\to Q_{k}$ in the Hausdorff distance. In other words, we need to show the following:
  1. (i)
    if $a\in Q_{k}$, then there exists $a_{n}\in Q^{n}_{k}$ such that $a_{n}\to a$,
  2. (ii)
    if $a_{n}\in Q^{n}_{k}$ and $a_{n}\to a$, then $a\in\overline{Q}_{k}$.
We prove these using the following technical claim.
Claim 3.3.
Let $(\mathcal{G}_{1},\mathcal{P}_{1}),(\mathcal{G}_{2},\mathcal{P}_{2}),\ldots$ be a sequence of edge-coloured graphings with at most $k$ colours that converges as coloured graphings to the coloured graphing $(\mathcal{G},\mathcal{P})$ in the local sense. Then, $\varphi_{n}/\mathcal{P}_{n}$ converges to $\varphi/\mathcal{P}$ in $\mathbb{R}^{2^{k}}$.
Proof of claim..
For $A\subseteq[k]$, let $\tilde{A}_{n}$ (respectively $\tilde{A}$) denote the set of edges of $\mathcal{G}_{n}$ (respectively $\mathcal{G}$) with colour in $A$. Clearly, $\varphi_{n}(\tilde{A}_{n})=(\varphi_{n}/\mathcal{P}_{n})(A)$, and similarly for $\varphi$, $\mathcal{P}$, and $\tilde{A}$. We denote by $\mathcal{H}_{A,n}$ the subgraphing of $\mathcal{G}_{n}$ with edge set $\tilde{A}_{n}$, and define $\mathcal{H}_{A}$ similarly. The coloured local convergence of the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$ implies that $\mathcal{H}_{A,n}$ locally converges to $\mathcal{H}_{A}$ for all $A$. As $f$ is continuous with respect to local convergence, we have
$(\varphi_{n}/\mathcal{P}_{n})(A)=\varphi_{n}(\tilde{A}_{n})=f(\mathcal{H}_{A,n})\to f(\mathcal{H}_{A})=\varphi(\tilde{A})=(\varphi/\mathcal{P})(A),$
concluding the proof of the claim. ∎
Proof of (i). As $a\in Q_{k}$, there is a Borel partition $\mathcal{P}$ of the edges of $\mathcal{G}$ such that $\varphi/\mathcal{P}=a$. By $\mathcal{P}$ being a Borel partition and by local-global convergence, we can find some Borel partition $\mathcal{P}_{n}$ of the edges of $\mathcal{G}_{n}$ such that the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$ locally converges to $(\mathcal{G},\mathcal{P})$ as coloured graphings [^1] Usually one defines local-global convergence using vertex colorings, but edge colorings can also be approximated, see [17, Proposition 19.15].. By Claim 3.3, the corresponding $a_{n}=\varphi_{n}/\mathcal{P}_{n}$ converge to $a$.
Proof of (ii). Assume $a_{n}\to a$. For all $a_{n}$, there exists a partition $\mathcal{P}_{n}$ such that $a_{n}=\varphi_{n}/\mathcal{P}_{n}$. By possibly taking a subsequence, we may assume that the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$ converges locally as coloured graphings. Let $(\mathcal{G}^{*},\mathcal{P}^{*})$ represent the limit. The graphing $\mathcal{G}^{*}$ a priori has nothing to do with $\mathcal{G}$ or the sequence $\mathcal{G}_{n}$ because local-global convergence of the sequence $\mathcal{G}_{n}$ does not imply a partition $\mathcal{P}$ of the edges of $\mathcal{G}$ such that $(\mathcal{G},\mathcal{P})$ represents the limit of $(\mathcal{G}_{n},\mathcal{P}_{n})$. Nevertheless, by completeness, such a coloured graphing $(\mathcal{G}^{*},\mathcal{P}^{*})$ exists. However, local-global convergence provides arbitrarily good approximations of $(\mathcal{G}^{*},\mathcal{P}^{*})$ on $\mathcal{G}$. That is, there exist Borel-partitions $\mathcal{R}_{n}$ of $\mathcal{G}$ such that the $(\mathcal{G},\mathcal{R}_{n})$ locally converge to $(\mathcal{G}^{*},\mathcal{P}^{*})$. By Claim 3.3 applied to the sequence $(\mathcal{G}_{n},\mathcal{P}_{n})$, we get that $a=\varphi_{\mathcal{G}^{*}}/\mathcal{P}^{*}$. Let $b_{n}=\varphi/\mathcal{R}_{n}$. Then, $b_{n}\in Q_{k}$ and applying the claim to $(\mathcal{G},\mathcal{R}_{n})$ gives $b_{n}\to a$ as well. Thus we get $a\in\overline{Q}_{k}$ as stated, finishing the proof of the theorem. ∎
With the help of Theorem 3.2, we are ready to prove Theorem 1.1.
Proof of Theorem 1.1..
Let $f(\mathcal{G})=\rho_{\mathcal{G}}(E(\mathcal{G}))$ be the graphing parameter that assigns to every graphing the total rank of its edge set. Then $\varphi_{\mathcal{G},f}=\rho_{\mathcal{G}}$, and Theorems 3.1 and 3.2 together complete the proof. ∎
3.2 Non-continuity for local convergence
Finally, we prove Theorem 1.2, showing that local convergence does not guarantee quotient-convergence of the matroids.
Proof of Theorem 1.2..
Let $(G_{1},G_{2},\dots)$ be a sequence of $3$-regular, $c$-expander graphs with $|V(G_{n})|=2n$, such that their girth $g_{n}$ tends to infinity. Let $\mathcal{G}$ be a subsequential local-global limit of $G_{n}$. Note that $\mathcal{G}$ is a 3-regular treeing. Let $H_{n}=G_{n}\cup G_{n}$, that is, $H_{n}$ is the disjoint union of two copies of $G_{n}$. Clearly, the sequence $H_{n}$ locally converges to $\mathcal{G}$ as well. Write $\rho_{G_{n}}$ for $r_{G_{n}}/|V(G_{n})|$, and similarly for $H_{n}$.
The partition of $H_{n}$ into two copies of $G_{n}$ defines a factor of the cycle matroid:
$a_{n}\in Q_{2}(H_{n})\subseteq\mathbb{R}^{2^{2}},\ a_{n}(\emptyset)=0,\ a_{n}(\{1\})=a_{n}(\{2\})=\frac{|V(G_{n})|-1}{2|V(G_{n})|},\ a_{n}(\{1,2\})=\frac{2|V(G_{n})|-2}{2|V(G_{n})|}.$
The limit of $a_{n}$ is
$a\in\mathbb{R}^{2^{2}},\ a(\emptyset)=0,\ a(\{1\})=a(\{2\})=\frac{1}{2},\ a(\{1,2\})=1.$
[…]
By the definition of expander sequences, for every $S\subseteq V(G_{n})$ with $|S|\leq n$, the number of edges connecting $S$ to $V(G_{n})\setminus S$ is at least $c\cdot|S|$. Let $E(G_{n})=E_{1}\cup E_{2}$ be any 2-partition of the edges. For $i\in\{0,1,2,3\}$, we denote by $V_{i}$ the set of vertices of $G_{n}$ with degree $i$ in $E_{1}$ and let $m_{i}=|V_{i}|$. Suppose that this partition results in a quotient closer to $a$ than $\varepsilon$. Then, $|\rho_{G_{n}}(E_{i})-\frac{1}{2}|<\varepsilon$ for $i=1,2$. Let us note that $r_{G_{n}}(E_{1})\leq 2n-m_{0}$, and the previous two inequalities imply $m_{0}/(2n)\leq(1/2+\varepsilon)$. The same bound holds for $m_{3}$.
We get a contradiction by showing that this is not possible if $g_{n}\geq 32/c$ and $\varepsilon<c/32$. We bound the number of $E_{1}$-components as follows. The number of singletons is clearly $m_{0}$. If an $E_{1}$-component is not a tree, then it has at least $g_{n}$ vertices, and hence the number of such components is at most $2n/g_{n}$. Finally, if a component is a non-singleton tree, then it has at least $2$ vertices off degree $1$, so the number of such components is bounded by $m_{1}/2$. We get that
$\rho_{G_{n}}(E_{1})\geq 1-\frac{1}{2n}\left(m_{0}+\frac{m_{1}}{2}+\frac{2n}{g_{n}}\right).$
Writing up the analogous inequality for $E_{2}$ and adding up the two inequalities, we get
$\rho_{G_{n}}(E_{1})+\rho_{G_{n}}(E_{2})\geq 2-\frac{1}{2n}\left(m_{0}+m_{3}+\frac{m_{1}+m_{2}}{2}\right)-\frac{2}{g_{n}}=1+\frac{m_{1}+m_{2}}{4n}-\frac{2}{g_{n}}.$ (3)
Without loss of generality, we may assume that $m_{0}\leq m_{3}$, implying $m_{0}\leq n$. The neighbours of $V_{0}$ in $G$ are contained in $V_{1}\cup V_{2}$, and hence
$m_{1}+m_{2}\geq cm_{0}.$
We also have
$m_{1}+m_{2}=2n-m_{0}-m_{3}\geq 2n-m_{0}-\left(\frac{1}{2}+\varepsilon\right)2n=\left(\frac{1}{2}-\varepsilon\right)2n-m_{0}.$
These two inequalities imply
$m_{1}+m_{2}\geq\frac{c}{c+1}\left(\frac{1}{2}-\varepsilon\right)2n.$
Thus, by (3) and the assumptions $g_{n}\geq 32/c$ and $\varepsilon<c/32$, we get
$\rho_{G_{n}}(E_{1})+\rho_{G_{n}}(E_{2})\geq 1+\frac{c}{2c+2}\left(\frac{1}{2}-\varepsilon\right)-\frac{2}{g_{n}}>1+2\varepsilon,$
contradicting $|\rho_{G_{n}}(E_{i})-\frac{1}{2}|<\varepsilon$ for $i=1,2$. ∎
4 Exposed minorizing measures and hyperfinite spanning forests
In this section, we focus on characterizing the exposed points of the base-polytope of the cycle matroid of a graphing. Recall that for a finite matroid $M=(E,r)$, the convex hull of its bases is described as
$C_{M}=\left\{x\in\mathbb{R}^{E}~{}\Big{|}~{}x\geq 0,\ \sum_{e\in A}x(a)\leq r(A)\text{ for }A\subseteq E,\ \sum_{e\in E}x(e)=r(E)\right\}.$
Let $\mathcal{G}=(X,E,\mu)$ be a graphing, let $\tilde{\mu}$ denote its edge measure, and $\rho$ the corresponding rank function. The analytic counterpart of $C_{M}$ is the convex set of basic minorizing finitely additive measures
$C_{\mathcal{G}}=bmm^{+}(\rho)=\big{\{}\alpha\textrm{ finitely additive measureon $E$}~{}\big{|}~{}\alpha\leq\rho,\ \alpha(E)=\rho(E)\big{\}}.$
In this section we write $C=C_{\mathcal{G}}$ for brevity. If $\mathcal{G}$ has no infinite components, then $\rho(E)=1$. Furthermore, as $\alpha\leq\rho\leq\tilde{\mu}$, in fact all $\alpha\in C$ are $\sigma$-additive measures and they admit Radon-Nikodym derivatives with respect to $\tilde{\mu}$, denoted by
$f_{\alpha}(x)=\frac{\partial\alpha}{\partial\tilde{\mu}}(x).$
Note that $f_{\alpha}$ takes values in $[0,1]$. If $f_{\alpha}$ is 00-$1$ valued $\tilde{\mu}$-almost everywhere, then $\alpha$ is clearly extreme in $C$. In this case there is a corresponding measurable set of edges $E_{\alpha}=\{e\in E\mid f_{\alpha}(e)=1\}\subseteq E$. The first observation we make is that such $\mathcal{T}=(X,E_{\alpha},\mu)$ is an essential hyperfinite spanning forest.
Lemma 4.1.
If $f_{\alpha}$ is 0-1 valued, then $\mathcal{T}=(X,E_{\alpha},\mu)$ is an essential hyperfinite spanning forest.
Proof.
By $\alpha\leq\rho$, $E_{\alpha}$ contains no cycles, so $\mathcal{T}$ is a forest. As $\alpha\in\text{bmm}^{+}(\rho_{\mathcal{G}})$, we have $\rho_{\mathcal{G}}(E_{\alpha})=\rho_{\mathcal{G}}(E)$, and therefore $\mathcal{T}$ is essential. Also, we have $\textrm{cost}(\mathcal{T})=\tilde{\mu}(E_{\alpha})=\alpha(E)=\rho_{\mathcal{T}}(E_{\alpha})$, so $\mathcal{T}$ is hyperfinite by Proposition 2.2. ∎
For an essential hyperfinite spanning forest $\mathcal{T}=(X,F,\mu)$, let $\alpha_{F}$ denote the measure defined by $\alpha_{F}(A)=\tilde{\mu}(A\cap F)$. We have the following.
Lemma 4.2.
For any essential hyperfinite spanning forest $\mathcal{T}=(X,F,\mu)$ of a graphing $\mathcal{G}$, $\alpha_{F}$ is in $C$.
Proof.
Let $A\subseteq E$ be a Borel subset of the edges of $\mathcal{G}$. Then $\mathcal{T}_{A}=(X,A\cap F,\mu)$ is a hyperfinite forest and $\tilde{\mu}(A\cap F)=\textrm{cost}(\mathcal{T}_{A})$. By Proposition 2.2, we get $\textrm{cost}(\mathcal{T}_{A})=\rho(A\cap F)$. By monotonicity of $\rho$, this implies $\alpha_{F}(A)\leq\rho(A)$. ∎
Remark 4.3.
Observe that in the above argument, the equality $\alpha_{F}(A)=\rho(A)$ holds if and only if $\rho(A\cap F)=\rho(A)$. This happens exactly when for $\mu$-almost all $x\in X$ the $A$- and $(A\cap F)$-components of $x$ have the same size. In case this size is infinity, the $(A\cap F)$-component does not need to be the whole $A$-component. In other words, there is equality exactly when $\mathcal{T}_{A}=(X,A\cap F,\mu)$ is essential in $\mathcal{G}_{A}=(X,A,\mu)$.
4.1 Exposed points
We say that $f_{\alpha}$ is exposed by $g\in L^{1}(E,\tilde{\mu})$ if the corresponding continuous linear functional $\varphi_{g}$ of $L^{\infty}(E,\tilde{\mu})$ attains its strict maximum on $C$ at $f_{\alpha}$. More formally, if for any $\beta\in C$, $\beta\neq\alpha$ we have $\varphi_{g}(f_{\alpha})>\varphi_{g}(f_{\beta})$. Clearly, exposed points are extreme. We prove that all exposed points of $C$ are of the form $\alpha_{F}$, or in other words, all the exposed $f_{\alpha}$ are 00-$1$ valued.
Theorem 4.4.
Let $f_{\alpha}\in C$ be exposed by a function $g\in L^{1}(E,\tilde{\mu})$. Then, $f_{\alpha}$ is 00-$1$ valued $\tilde{\mu}$-almost everywhere.
Remark 4.5.
Usually, for a convex set $C$ in a topological vector space, a point $x\in C$ is called exposed if there exists a continuous linear functional that attains its strict maximum at $x$. In our case, $C\subseteq L^{p}(E,\tilde{\mu})$ for all $p\in[1,\infty]$. For $f_{\alpha}$ to be exposed in $L^{p}(E,\tilde{\mu})$ where $p\neq\infty$, the witness $g$ has to be in $L^{q}(E,\tilde{\mu})$, where $q$ is the dual pair of $p$, that is, $11/p+1/q=$ 1. This means that it is easier to be exposed for higher values of $p$ since there are more functions in $L^{q}$ for lower values of $q$. It is even easier to be exposed in $L^{\infty}(E,\tilde{\mu})$, even with the restriction that the linear functional corresponds to an $L^{1}$ function, and this is enough for our argument. What is not enough is to be exposed in $L^{\infty}(E,\tilde{\mu})$ in general, because the linear functional might correspond to a finitely additive measure, and not to a measure whose Radon-Nikodym derivative would be the witness $g\in L^{1}(E,\tilde{\mu})$.
Proof of Theorem 4.4..
We treat $g$ as a weight function on the edges, and our goal is to find a maximal essential hyperfinite spanning forest $F$ with respect to this weight function. We write $\alpha_{F}$ for the corresponding measure in $C$, i.e. $\alpha_{F}(A)=\tilde{\mu}(A\cap F)$ for all measurable $A\subseteq E$. The algorithm will result in an $F$ satisfying $\alpha_{F}(\{g\geq t\})=\rho(\{g\geq t\})$. In a layer-cake integration this implies, for all $\beta\in C$,
[…]
We run a version of Kruskal’s greedy algorithm [14] on $\mathcal{G}$. To handle ties, we fix an arbitrary Borel ordering of $E$ in advance, and whenever there is a tie in the $g$ values among finitely many edges considered, we choose the maximal one according to the Borel ordering. The algorithm constructs Borel edge sets $F_{1}\subseteq F_{2}\subseteq\ldots$, with $F=\bigcup F_{n}$ being the final output. In the first step, every vertex simultaneously chooses the incident edge with the highest $g$-value to be included in $F_{1}$. During the $n^{\mathrm{th}}$ step, every finite component of $F_{n-1}$ adds the edge with the highest $g$-value from the set of edges leaving it, and thus we obtain $F_{n}$. In case there is no edge leaving a finite $F_{n-1}$ component we do not add anything, but observe that in this component $F$ is already a spanning tree. Also note that infinite components do not add any edges, but it may happen that a finite component adds an edge that connects it to an infinite one.
It is not difficult to see that the sets $F_{n}$ are Borel, cycles cannot be created, and $F$ is essential. We claim that $\tilde{\mu}(F)=\rho_{\mathcal{G}}(E)$, so $F$ is hyperfinite by Proposition 2.2. Indeed, set up a payment function as follows. In the beginning, every vertex receives $1$ token. In a general step, every finite connected component gives $1$ token to the edge it chooses. After the addition of the edges, for every finite connected component there is exactly one edge who received $2$ tokens. This edge gives $1$ token back to the component, thus ensuring that every finite component has a token before the next step. Should a finite $F$-component become spanning in a finite $\mathcal{G}$-component, and therefore has no use for its token, this token is redistributed equally among the vertices of this component. In the end, vertices in finite $\mathcal{G}$-components paid $1-\frac{1}{|\mathcal{G}_{x}|}$ token, and vertices in infinite $\mathcal{G}$-components paid 1 token. On the other hand every chosen edge received 1 token, hence $\rho_{\mathcal{G}}(E)=\tilde{\mu}(F)$ by the Mass Transport Principle. As $F$ is hyperfinite, by Lemma 4.2, we have $\alpha_{F}\in C$.
Let $A_{t}=\{g\geq t\}$ and $F_{\geq t}=F\cap A_{t}$. We further observe that $F_{t}$ is an essential spanning forest of $A_{t}$. That is, for $\mu$-almost every $x\in X$,
  1. (1)
    if the $A_{t}$-component of $x$ is finite, then $F_{\geq t}$ is a spanning tree inside this component, and
  2. (2)
    if the $A_{t}$-component of $x$ is infinite, then its $F_{\geq t}$-component is also infinite – though it might be smaller.
For (1), note that the edges leaving the $A_{t}$-component all have $g$-values strictly less than $t$, while being an $A_{t}$-component means that it can be connected by edges with $g$-value at least $t$. So the algorithm will connect these points before growing the component any further. For (2), the reasoning is analogous. The conclusion, however, is different, This is due to the fact that we cannot guarantee the connectedness of $F_{\geq t}$ inside the $A_{t}$-component, because the $F_{\geq t}$-components stop growing once they are all infinite.
By Remark 4.3, (1) and (2) imply $\alpha_{F}(A_{t})=\rho(A_{t})$, completing the proof of the theorem. ∎
Let us make a couple of remarks.
Remark 4.6.
If $f_{\alpha}$ is 00-$1$ valued, then it is exposed – the witness for this is simply choosing $g=f_{\alpha}$. In fact, in this case $f_{\alpha}$ is strongly exposed, i.e. for any sequence $\beta_{n}\in C$ we have $\varphi_{g}(f_{\beta_{n}})\to\varphi_{g}(f_{\alpha})\Rightarrow||f_{\beta_{n}}-f_{\alpha}||_{p}\to 0$ for all $p\in[1,\infty)$. So Theorem 4.4 characterizes exposed points of $C$.
Remark 4.7.
By the Klee–Straszewicz theorem [13, Theorem 2.1], exposed points are dense among extreme points of $C$ in the weak topology of $L^{\infty}(E,\tilde{\mu})$. Moreover, the exposed points determine $C$ as $C=\overline{\textrm{conv}(\exp(C))}^{w^{*}}$.
4.2 Weakly exposed points
Even in finite dimension, extreme points are not necessarily exposed. To see this, consider the following set in $\mathbb{R}^{2}$:
$S=\big{\{}(x,y)\mid-1\leq x\leq 1,\ -1\leq y\leq 1\big{\}}\cup\big{\{}(x,y)\mid x^{2}+y^{2}\leq 1\big{\}}.$
Then $(-1,0)$ and $(1,0)$ are clearly extreme points. However, there is a unique supporting hyperplane through any of them: $x=-1$ through $(-1,0)$ and $x=1$ through $(1,0)$, hence these points are not exposed. Nevertheless, these extreme points can be exposed in finitely many steps by iteratively restricting the problem to the intersection of the convex set and a supporting hyperplane. In this subsection, we use this idea to generalize Theorem 4.4 to extreme points that we call weakly exposed.
We say that an extreme point $x\in C$ is exposed in two steps by $L^{1}$ functions, if one can find $g_{1},g_{2}\in L^{1}(E,\tilde{\mu})$ such that $\varphi_{g_{1}}(x)$ is maximal in $C$, and $x$ to strictly maximizes $\varphi_{g_{2}}$ in $C^{\prime}=\{y\in C\mid\varphi_{g_{1}}(y)=\varphi_{g_{1}}(x)\}$. Note that the proof of Theorem 4.4 yields the same conclusion for points that are exposed in two steps. The only difference is that during the algorithm, we always choose an edge that is maximal with respect to $g_{1}$, and in case of ties, we choose the one maximal with respect to $g_{2}$. Finally, if there is still a tie, we decide according to the Borel ordering we fixed at the beginning.
More generally, let $\alpha\in\omega_{1}$ be a countable ordinal and let $\mathbf{g}\colon\alpha\to L^{1}(E,\tilde{\mu})$ be a list of functions. Let us define $\mathbf{g}(x)\in\mathbb{R}^{\alpha}$ by $f(x)_{\beta}=\mathbf{g}(\beta)(x)$. Then, $\mathbf{g}$ defines a Borel partial ordering on $C$ based on the lexicographical ordering of the images of points of $C$ in $\mathbb{R}^{\alpha}$. We denote this ordering by $<_{\mathbf{g}}$ and call a point $x\in C$ exposed in $\alpha$ steps by the list $\mathbf{g}$ if $x$ is the unique maxima of $<_{\mathbf{g}}$ in $C$. In other words, for all $y\in C$ there exists an ordinal $\beta_{y}\leq\alpha$ such that for all $\gamma\leq\beta_{y}$, we have $\mathbf{g}(\gamma)(x)=\mathbf{g}(\gamma)(y)$ and $\mathbf{g}(\beta_{y})(x)>\mathbf{g}(\beta_{y})(y)$. We say $x\in C$ is weakly exposed, if it is exposed in $\alpha$ steps for some countable ordinal $\alpha$.
Theorem 4.8.
Let $f_{\alpha}\in C$ be weakly exposed by a list of functions in $L^{1}(E,\tilde{\mu})$. Then, $f_{\alpha}$ is 00-$1$ valued $\tilde{\mu}$-almost everywhere.
Proof.
The proof of Theorem 4.4 holds for weakly exposed points as well by always choosing the $<_{\mathbf{g}}$-maximal edge. ∎
Corollary 4.9.
Weakly exposed points of $C$ are strongly exposed.
In finite dimension, all extreme points of a compact convex set are weakly exposed. In infinite dimension, however, this is not necessarily true. The closed unit ball of $\ell^{p}(\mathbb{N})$ for $21<p<$ 2 in $\ell^{2}(\mathbb{N})$ has extreme points that are not even support points, see [23, Example 7.5].
Remark 4.10.
The algorithm in the proof of Theorem 4.4 uses an arbitrary Borel ordering $<$ on the edge set. The resulting forest is called the Wired maximal spanning forest, see  [20, Chapter 11]. Usually, such forests are studied with the ordering coming from iid $[0,1]$ labels on the edges of a countable graph.
5 Duality
In this section, we present various observations based on the idea that in an infinite planar graph, the complement of an essential spanning forest is an essential spanning forest of the dual.
Roughly speaking, a planar graphing is a graphing $\mathcal{G}=(X,E,\mu)$ whose connected components are planar and are also endowed with a concrete embedding in a measurable way. The embedding is encoded by specifying the finite facial cycles of the components, and we assume this collection of edge sets to be Borel inside the standard Borel space of finite subsets of $E$. For the precise definition, see [7, Section 3.2]. The component of a $\mu$-random vertex, together with the embedding is a unimodular random map in the sense of [1].
We will, in fact, restrict our attention to the case when all facial cycles of $\mathcal{G}$ are finite, and refer to them as faces. The assumption of $2$-vertex-connectivity and bounded size of faces mentioned in the Introduction also applies. The fact that the encoding is Borel ensures that we can talk about the dual graphing $\mathcal{G}^{*}=(X^{*},E^{*},\mu^{*})$, where $X^{*}$ is the set of faces of $\mathcal{G}$. There is a measure preserving Borel bijection $\sigma\colon(E,\tilde{\mu})\to(E^{*},\tilde{mu}^{*})$ that denotes the usual correspondence: $\sigma(e)$ connects two distinct faces $x_{1}^{*}$ and $x_{2}^{*}$ if and only if $e$ is a common boundary edge of $x_{1}^{*}$ and $x_{2}^{*}$. For sake of simplicity, we write $e^{*}=\sigma(e)$ and $F^{*}=\sigma(F)$ for $F\subseteq E(\mathcal{G})$. For an edge set $F\subseteq E$, we use the notation $F^{*}=\{e^{*}\mid e\in F\}$.
Examples of planar graphings include Schreier graphings of p.m.p. actions of finitely generated groups with planar Cayley graphs (in particular surface groups), as well as Palm graphings of planar factor graphs of free invariant point processes on $\mathbb{R}^{2}$ or $\mathbb{H}^{2}$ [21, Remark 17].
Remark 5.1.
For graphings, we usually assume $\mu(X)=1$ for convenience. For finite graphs this corresponds to normalization by the number of vertices. However, when duals are considered, this normalization is less natural, since the dual graph typically does not have the same number of vertices. It would probably be more natural to normalize by the number of edges which is the same for the two graphs. Nevertheless, to be consistent with the rest of the paper, we do not change the normalization, and simply accept that the vertex measure of the dual graphing is not necessarily 1.
Remark 5.2.
Limits of sequences of finite planar graphs are hyperfinite by the Lipton-Tarjan Planar Separation Theorem [16]. Nevertheless, planar graphings need not be hyperfinite. In fact, there is a sharp dichotomy between hyperfinite and non-hyperfinite planar graphings, see [1] and [26].
We define the cocyclic matroid of $\mathcal{G}$ by the infinite analogue of the usual rank function formula:
$\rho^{D}_{\mathcal{G}}(A)=\rho_{\mathcal{G}}(E\setminus A)+\tilde{\mu}(A)-\rho_{\mathcal{G}}(E).$
Note that $\rho_{\mathcal{G}}(E)=1$, and $\rho^{D}_{\mathcal{G}}$ is submodular as it is the sum of a submodular and a modular function. Moreover, the rank of the full edge set in the cocycle matroid is $\rho^{D}_{\mathcal{G}}(E)=\tilde{\mu}(E)-\mu(X)$. There is a nice correspondence between the elements of the associated base-polytopes:
$\alpha\in\text{bmm}^{+}(\rho_{\mathcal{G}})\iff\tilde{\mu}-\alpha\in\text{bmm}^{+}(\rho^{D}_{\mathcal{G}}).$
5.1 Duality for hyperfinite planar graphings
For finite planar graphs, the cocycle matroid is the cycle matroid of the dual. In the Borel setting this holds only for hyperfinite planar graphings. In fact, we will exploit exactly the discrepancy between $\rho_{\mathcal{G}}$ and $\rho^{D}_{\mathcal{G}^{*}}$ in Subsection 5.2 to prove Theorem 1.9.
Remark 5.3.
Theorem 1.8 is actually a reformulation of Euler’s formula, stating that for any embedding of a finite planar graph, the equality $n-m+f=2$ holds, where $n$, $m$ and $f$ denote the number of vertices, edges and faces, respectively. For an planar graphing $\mathcal{G}$ with infinite components, the special case $\rho^{D}_{\mathcal{G}*}(E)=\rho_{\mathcal{G}}(E)$ translates into $\tilde{\mu}(E)-\mu^{*}(X^{*})=\mu(X)$, that is, the $2$ in the formula disappears in the limit.
We present two proofs for Theorem 1.8, one relying on Theorem 4.4, the other using the Mass Transport Principle.
Proof of Theorem 1.8 using Theorem 4.4.
We noted earlier that $\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})=\tilde{\mu}-\text{bmm}^{+}(\rho_{\mathcal{G}^{*}})$. By Theorem 4.4, we know that exposed points of $\text{bmm}^{+}(\rho_{\mathcal{G}})$ and $\text{bmm}^{+}(\rho_{\mathcal{G}^{*}})$ are essential hyperfinite spanning forests of $\mathcal{G}$ and $\mathcal{G}^{*}$, respectively. Note also that they are in a one-to-one correspondence by complementation. Here we use that $\mathcal{G}$ and $\mathcal{G}^{*}$ are hyperfinite, to ensure that the complement is again hyperfinite.
These observations together show that $\text{bmm}^{+}(\rho_{\mathcal{G}})$ and $\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$ have the same set of exposed points. This implies $\text{bmm}^{+}(\rho_{\mathcal{G}})=\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$, as they are the weak* closure of the convex hull of the same set, see Remark 4.7. Finally, this implies $\rho_{\mathcal{G}}=\rho^{D}_{\mathcal{G}^{*}}$ by the fact that a submodular function $\varphi$ is determined by $\text{bmm}^{+}(\varphi)$ in the sense that $\varphi(A)=\sup_{\alpha\in\text{bmm}^{+}(\varphi)}\alpha(A)$, see [19, Corollary 6.10]. The proof of $\rho^{D}_{\mathcal{G}}=\rho_{\mathcal{G}^{*}}$ is analogous. ∎
Proof of Theorem 1.8 using the Mass Transport Principle.
By using the definitions of $\rho_{\mathcal{G}}$ and $\rho^{D}_{\mathcal{G}^{*}}$, and the equalities $\tilde{\mu}^{*}=\tilde{\mu}$ and $\rho_{\mathcal{G}^{*}}(E^{*})=\mu^{*}(X^{*})$, we obtain that $\rho_{\mathcal{G}}(A)=\rho_{\mathcal{G}^{*}}(A)$ is equivalent to
[…]
To verify the equality, we set up a payment between vertices, faces, and edges. Motivated by the second integral, every face $x^{*}$ that is in a finite $(E^{*}\setminus A^{*})$-component pays $1/|V(\mathcal{G}^{*,E^{*}\setminus A^{*}}_{x^{*}})|$ token to the component. This way every finite $(E^{*}\setminus A^{*})$-component collects 1 token in total, and this token is passed on to an edge (chosen in a Borel way) in the $A$-cycle surrounding the component. We can also make sure that the edges that get payed, oriented from the component that pays to them, form an acyclic oriented graph on the components of $E^{*}\setminus A^{*}$. This way the edges of $A$ that do not receive a payment form a subforest $A^{\prime}\subseteq A$ that has the same connected components in $\mathcal{G}$ as $A$.
Next, every $x\in X$ in a finite $A$-component pays $1/|V(\mathcal{G}^{A}_{x})|$ token to every $A^{\prime}$-edge in its component. This way all such vertices pay $1-1/|V(\mathcal{G}^{A}_{x})|$, and such $A$-edges receive $1$ token in total.
What remains are infinite $A^{\prime}$ components and their vertices. Write $A^{\prime\prime}$ for this edge set, and $X^{\prime\prime}$ for the vertices. Then $\tilde{\mu}(A^{\prime\prime})=\mu(X^{\prime\prime})$ by Proposition 2.2, as $A^{\prime\prime}$ is a hyperfinite treeing on $X^{\prime\prime}$ with infinite components. In a mass transport formulation, this means that we can arrange that every vertex in $X^{\prime\prime}$ pays 1 token, and every edge in $A^{\prime\prime}$ gets 1 token.
Concluding the above, vertices and faces payed in total the right hand side of (4), while every edge in $A$ received 1, so (4) holds by the Mass Transport Principle. ∎
5.2 A matroid attaining the cost
Proof of Theorem 1.9.
As we are in the non-hyperfinite case, the dual $\mathcal{G}^{*}$ has no $2$-ended components $\mu^{*}$-almost surely. By [7], one can pick a 1-ended hyperfinite spanning forest $F^{*}\subseteq\mathcal{G}^{*}$. The corresponding measure $\alpha_{F^{*}}$ is in $\text{bmm}^{+}(\rho_{\mathcal{G}^{*}})$, implying $\tilde{\mu}-\alpha_{F^{*}}\in\text{bmm}^{+}(\rho^{D}_{\mathcal{G}^{*}})$. This measure, considered on $E(\mathcal{G})$, corresponds to $E\setminus F$. The key idea, also exploited in [7], is that $F^{*}$ being $1$-ended almost everywhere means that $E\setminus F$ is connected on almost every component of $\mathcal{G}$. That is, $E\setminus F$ is a treeing of $\mathcal{G}$. Therefore, $\rho^{D}_{\mathcal{G}^{*}}(E)=\alpha_{E\setminus F}(E)=\tilde{\mu}(E\setminus F)=\textrm{cost}(\mathcal{G})$, where the last inequality holds by Theorem 2.1. ∎
5.3 Free versus wired spanning forests
Let $\mathcal{G}$ be a graphing with a Borel order $<$ on the edges. We have already seen in Remark 4.10 that the algorithm for finding bases with respect to $\rho_{\mathcal{G}}$ produces the wired maximal spanning forest, denoted by $\mathcal{T}_{w}$. There is a natural counterpart, the free maximal spanning forest. We again refer to [20, Chapter 11] for a thorough overview, with the caveat that here we treat arbitrary Borel orders on the edge set.
We define the free maximal spanning forest $\mathcal{T}_{f}$ by simultaneously deleting all edges $e$ of $\mathcal{G}$ for which there exists a cycle $C$ containing $e$ in which $e$ is $<$-minimal. It is straightforward to check that $\mathcal{T}_{f}$ is acyclic, and all of its components are infinite. Moreover, $\mathcal{T}_{w}\subseteq\mathcal{T}_{f}$.
Proof of Theorem 1.10.
When $\mathcal{G}$ and $\mathcal{G}^{*}$ are dual planar graphings and the Borel total ordering is the same on $E$ and $E^{*}$, the free maximal spanning forest of $\mathcal{G}$ is exactly the complement of the wired minimal spanning forest of $\mathcal{G}^{*}$. ∎
Remark 5.4.
Based on Theorem 1.10, one could hope to use the free maximal spanning forest for arbitrary, not necessarily planar graphings to define an interesting matroid that goes beyond hyperfinite spanning forests. The problem is that the free maximal spanning forest built on two different Borel orders might have different edge measure, and so they cannot form bases of the same matroid. For example, let $\mathcal{G}$ be the Schreier graphing of a free action of $F_{2}\times\mathbb{Z}$. Let $<_{1}$ be such that all $F_{2}$-edges are larger than all $\mathbb{Z}$-edges, and the ordering is reversed for $<_{2}$. Then, $\mathcal{T}^{<_{1}}_{f}$ contains exactly the $F_{2}$-edges while $\mathcal{T}^{<_{2}}_{f}$ contains exactly the $\mathbb{Z}$-edges, so their edge measures are $2$ and $1$, respectively.
Acknowledgement.
The authors are grateful to Miklós Abért, Boglárka Gehér, András Imolay, Gábor Kun, Balázs Maga, Nicolas Monod, Gábor Pete, Carsten Thomassen, Tamás Titkos and Dániel Virosztek for helpful discussions.
Márton Borbényi was supported by the ÚNKP-23-3 New National Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund (NKFIH). László Márton Tóth was supported by the NKFIH grant KKP-139502, ”Groups and graph limits”. The research was supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, and by Dynasnet European Research Council Synergy project – grant number ERC-2018-SYG 810115.